OpenStudy (chrisplusian):

use the shell method to find the volume of the solid generated by revolving the region bounded by the given curves about the given lines. y=9-x^2, y=9,x=3 revolve around the line y=9. Can someone check to see if I set the problem up correctly? I can do the integration.... I think

OpenStudy (chrisplusian):

\[2\pi \int\limits_{0}^{9}(3-\sqrt{9-y})(y) dy\]

OpenStudy (chrisplusian):

@Algebraic!

OpenStudy (noelgreco):

That's the integral!

OpenStudy (chrisplusian):

thank you

OpenStudy (chrisplusian):

do you concur @Algebraic!

OpenStudy (noelgreco):

Sure. That IS a 9 you have at the upper limit of integration, right?

OpenStudy (chrisplusian):

Yes

OpenStudy (anonymous):